OpenStudy (anonymous):

How to solve? Help please! Integral problem below(: ----The integral must be evaluated by breaking it up into a sum of three integrals.

OpenStudy (anonymous):

The integral \[\int\limits_{-1}^{3} |7x^2-x^3-6x| dx\] is split into three separate sums: \[\int\limits_{-1}^{0} |7x^2-x^3-6x| dx + \int\limits_{0}^{c} |7x^2-x^3-6x| dx +\int\limits_{c}^{3} |7x^2-x^3-6x| dx \] --->Find c. The values of each integral and the sum of all three.

OpenStudy (anonymous):

I figured out c. haha it's 1.

OpenStudy (anonymous):

well \(-x^3+7x^2-6x=-x(x^2-7x+6)=-x(x-1)(x-6)\)

OpenStudy (anonymous):

hence our discontinuities are at \(x=0,x=1,x=6\) as our absolute value changes "direction"

OpenStudy (anonymous):

yes, \(c=1\)

OpenStudy (anonymous):

so now i just evaluate each integral right?

OpenStudy (anonymous):

you gotta figure out over what intervals it is positive, and over what intervals it is negative if it is positive, then \(|f(x)|=f(x)\) and if it is negative, then \(|f(x)|=-f(x)\)

OpenStudy (anonymous):

but because in with in |abs| doesn't the function always come out positive? @satellite73

OpenStudy (anonymous):

@pdd21 you don't know a way of integrating arbitrary absolute values, though, so you need to determine (like @satellite73 said) where the function is positive and negative so you know how to simplify each integrand over its respective interval

OpenStudy (anonymous):

$$-x^3+7x^2-6x=-x(x^2-7x+6)=-x(x-1)(x-6)$$so we know how it factorizes... observe since the leading coefficient is negative and in the long run it dominates our behavior is a lot like \(-x^3\) i.e. its shape looks like this: